# thanks: Corresponding author

### v1

@@ I Introduction
Neutral atoms, such as 85Rb and 88Sr, often encode the qubit computational basis states in the electronic ground state manifold or clock states. These states have long coherence times on the order of seconds Jenkins et al. (2022); Ma et al. (2022) and the clock transition has a narrow linewidth Norcia et al. (2018). In the rest of this Article, we work in the $\{|0\rangle,|1\rangle,|r\rangle\}$ qutrit manifold of the 88Sr atom, which is endowed with the auxiliary Rydberg state $|r\rangle$ that facilitates long-range entanglement, which we leave unspecified. Fig. 1 shows the level scheme of 88Sr, and in which electronic states the $|0\rangle$- and $|1\rangle$-states have been embedded. The transition from $|1\rangle$ to $|r\rangle$ is driven by a single-photon process. Though we focus on ${}^{88}\text{Sr}$, our method applies to all neutral atoms with a similar level structure.
@@ II.1 Error modelling
Given is a noiseless unitary evolution $\mathcal{F}(\rho)=U\rho U^{\dagger}$. Introducing a noise channel with multi-qudit Kraus representatives $\{\mathfrak{D}_{\mu}\}$, we obtain an error channel
@@
relative to the perfect coherent channel $\mathcal{F}(\rho)$, for some coefficients $[\lambda_{\mu\nu}]$. We can transform this form to the Pauli basis as
@@
and use the Pauli twirling approximation Geller and Zhou (2013); Dankert et al. (2009); Jandura and Pupillo (2024) to obtain
@@
The validity of twirling is discussed in Appendix A. The Kraus operators satisfy the normalisation condition
on the Hilbert space $\mathfrak{H}^{m}$ of $m$ qudits. The operators themselves are generated by single-qudit Kraus operators $E_{q_{i}}^{i}$ acting on qubit $i$, with index $q_{i}$.
@@ II.2 Generating entanglement
Let $\mathbb{S}=\mathbb{S}_{X}\cup\mathbb{S}_{Z}$ be the group of stabilisers. Stabilisation between data qubits and stabiliser qubits $S\in\mathbb{S}$ is achieved through local 2-qubit entanglement operations. Neutral atoms provide a multitude of implementations, such as the simple constant pulses proposed by Jaksch et al. Jaksch et al. (2000), a time-optimal pulse proposed by Jandura et al. Jandura and Pupillo (2022), or a pulse that is robust against time-dependent variations in the laser control parameters proposed by Mohan et al. Mohan et al. (2023). Fig. 3(a) provides an implementation scheme of 5-qubit entanglement on the surface code. We implement subsequent 2-qubit gates in a clockwise fashion, and assume no interaction between Rydberg-excited data qubits. Movable tweezers, depicted in Fig. 3(b), can transport atoms within their stabilisers’ Rydberg blockade regime, where controlled logic can be performed, before moving them back to their initial positions in the lattice. Fig. 3(c) shows a complete schematic overview of how quantum error correction can be achieved on a neutral atom platform, showing an architecture with different dedicated zones for data storage, entanglement and readout.
@@
Figure 2: (a) The $[[9,1,3]]_{2}$ rotated surface code, where black circles denote data qubits that carry the 2 logical degrees of freedom, blue faces denote $X$-stabilisers and golden faces denote $Z$-stabilisers. (b) For each stabiliser type, the relevant circuit is shown for 2-plaquettes. $H$ is the Hadamard gate and meas denotes a mid-circuit measurement. 4-plaquettes have analogous circuit designs.
@@
Figure 3: (a) Entanglement protocol for 4-plaquettes. The dashed circle indicates the Rydberg blockade radius, limited below the interatomic distance. Controlled logic between stabiliser and data qubits is performed by transporting the neighbouring qubits within the blockade radius, and applying an appropriate series of pulses before moving it back to its original position in the lattice. We subsequently entangle other data qubits in a clockwise fashion, as indicated by the grey arrows. (b) Visualisation of atom transport, with all relevant timescales indicated below Bluvstein et al. (2022); Tian et al. (2023). AOM (acousto-optical modulator) tweezers have the ability to pick up an atom from its stationary trap and move it to a different location. Such transport allows for flexibility during execution of an algorithm, though transpilation may pose a serious challenge Tan et al. (2024). (c) Schematic architecture for neutral atom quantum computing experiments, tailored here for quantum error correction. We embed the CSS code in a 2D plane of atoms (grey zone), here seen from the side. Single-qubit operations are performed in the grey zone, entanglement is mediated in the purple Rydberg zone, and measurements of the qubit states are performed in the green readout zone. AOMs transport atoms between layers. The Rydberg laser that mediates the transition $|1\rangle\leftrightarrow|r\rangle$ is global, as well as the laser that reads out qubit states through fluorescence. Hence, dedicated zones are required to avoid crosstalk. Single-qubit operations in the grey zone are local and can be tuned from lattice site to site, within realistic experimental bounds.
@@ III Statistical model of errors
Figure 4: (a) A CSS code is mapped to a spin gauge model with random bonds, with spins $s^{X}$ (blue) and $s^{Z}$ (golden) representing stabilisers and bonds representing data qubits. Bonds between spins $s^{i}$ are ferromagnetic ($+J$) and favour long-range order, if the corresponding data qubit does not have an error that can be picked up by the stabiliser type $i$. Such bonds are depicted as blue dashed lines. (b) If a data qubit undergoes an error (red), triggering two flags (pink), its respective bond turns locally anti-ferromagnetic $(-J)$, indicated by a red dashed line.
@@ III.2 $\mathbb{Z}_{2}$ Gauge theory - Random plaquette model
Figure 5: (Top) Example of percolation on a 15$\times$15 spin system that is dual to the toric code, where spins are given by grey nodes. Bonds and vertices are randomly removed at a uniform rate $r=0.15$, leaving a subgraph on which top-bottom and left-right percolation can be found. Examples of logical support are indicated in green and blue respectively. As periodic boundary conditions apply, percolation clusters must connect left-right and top-bottom on the same horizontal respectively vertical line. (Bottom) Percolation probability as a function of the fraction of unerased qubits, for system sizes $L=5,10,25$. Darker colours indicate larger lattices. The vertical black line indicates the threshold at around 0.75 ($r\approx 0.25$.
@@ IV Code performance analysis
The rest of this Section is dedicated to obtaining qubit fidelities for the toric code implemented on a neutral atom quantum device, assuming two major loss channels mediated by the fragile Rydberg state. We evaluate the local error channels for 4-plaquettes, extracting probabilities $\chi_{\mu\mu}$ for certain twirled Pauli error strings on the data qubits, which includes the probability $q$ of a false flag error on the stabilisers, and insert these into a 3D random plaquette gauge model. Additionally, erasure rates are incorporated, whose effects are a result of the integrated Rydberg lifetime of the qubits plus the finite lifetime of the tweezer traps. Because of the anisotropy of our system, spacelike and timelike Wilson loops have different sensitivities to variations in temperature. Our calculations have shown that disorder is more spacelike than timelike, so that we sample timelike Wilson loops only. The details of the Metropolis algorithm and handling phase transitions are presented in Appendix C.
@@ IV.2 Handling erasure
Figure 9: (a) The phase diagrams for our quantum error correction implementation of the toric code, with stabiliser circuits being implemented by either the Jaksch or time-optimal protocol. The dark blue regions indicate the QEC regime where arbitrarily low logical qubit fidelities can be achieved, granted $d$ is sufficiently large. The light blue regions indicate regions where QEC is only viable if lost stabilisers are replenished from some fresh reservoir, alleviating some of the fidelity bounds. The blank regime entails all parameters $(\gamma,\bar{r})$ such that QEC is no longer possible. Error bars are standard Monte Carlo standard deviations for 2500 parallel runs. Because Wilson loop calculations for $r\gtrapprox 0.35$ are very difficult to perform in a Monte Carlo setting, we use a fit to approximate the phase boundaries, using the knowledge that for the limiting case of erasure-only errors ($\gamma\to 0$), we must recover 2D percolation thresholds calculated in Sec. III.3. A set of realistic system parameters are highlighted in black. (b) Typical graph of Wilson loop averages, normalised by perimeter/area. Each line is the average of 25 randomly chosen runs at $r=0$ and $\gamma/\Omega=10^{-3}$, for both the ordered ($T<T_{c}$) and disordered ($T>T_{c}$) regime, displaying a clear transition from a perimeter to an area law. The left graph shows $-\ln{\mathcal{W}[\tau]/|\tau|}$, while the right shows $-\ln{\mathcal{W}[\tau]/S_{\tau}}$. (c) Examples of when Wilson loops can be easily constructed. If the loop encloses no erased qubits, the enclosed area $S_{\tau}$ is equal to a Euclidean area. If some erasures are enclosed, then the area is effectively less. For larger $\bar{r}$, the calculation of the area/perimeter law for arbitrary loops becomes convoluted and imprecise.
@@ IV.3 Logical qubit fidelity and lifetime
$\gamma^{\text{Jaksch}}_{\text{th}}\approx 2.5\cdot 10^{-3}\Omega\quad\text{and}\quad\gamma^{\text{TO}}_{\text{th}}\approx 4.5\cdot 10^{-3}\Omega,$ (42)
which are consistent with literature on optimising neutral atom laser protocols for stabilisation for QEC Jandura and Pupillo (2024), as it provides an upper bound on error rate thresholds without requiring an (optimal) decoder. Note that the results agree very well because of the simplicity of the toric code, having local connectivity, and is compatible with the range of state-of-the-art experimental parameters
@@ Appendix A Twirling
For the Pauli twirling approximation, this corresponds to a physical insertion of Pauli operators, drawn at random on random locations within the circuit. It is not known exactly how accurate the twirling approximation is Geller and Zhou (2013); López et al. (2010), but it has been employed often with excellent agreement with exact simulations of the full system Ghosh et al. (2012). It is known that Pauli twirling reduces the density matrix to the space of 1-designs. Another option would be to invoke Clifford twirling, which forms a 3-design. However, the physical insertions of unitaries from the Clifford group $\mathcal{C}^{(2)}$ would cause our error propagation model to break down. These twirls are useful in efficient simulations nonetheless since it has been shown that random circuits with local connectivity are approximate $t$-designs and can be simulated classically using only polynomial time Brandão et al. (2016).
@@ Appendix B Calculating the entries of $[\chi_{\mu\nu}]$
On neutral atom quantum computers, measurement time is still the dominant timescale of a quantum error correction experiment, with $\tau_{\text{meas}}\approx 10$ ms being the order of magnitude for non-destructive fluorescence readout Covey et al. (2019). Therefore, we can assume that during measurement, all remaining Rydberg states will have decayed back to the $|1\rangle$-states (or erased). Choosing $\bar{\delta t}\gg 1/\gamma$, and $\bar{\mathfrak{D}}_{\mu}=\mathfrak{D}_{\mu}(\bar{\delta t})$, the final $\Gamma$-matrix becomes
@@ Appendix C Monte Carlo details
where $\varphi(\cdot)$ is some universal dimensionless function, and $\nu$ is a critical exponent Caracciolo et al. (1995); Palassini and Caracciolo (1999). Simulations were all performed on the GPU to boost efficiency, batching parallel runs by temperature points and error configurations. The plaquette gauge model has local gauge invariance with a local order parameter, so we can not employ FSS.
@@
We construct Wilson loops using square loops along the $xy$-planes of the model, since Pauli twirling shows there is more disorder along spacelike plaquettes than timelike plaquettes. We average over all Wilson loops that can be constructed, because of the isotropy of our model along spacelike plaquettes. In case of erasure, we only add loops to the thermodynamic averaging contribution that have no erased qubits on the loop, only inside. For $r\gtrapprox 0.15$, this signifies a major breakdown in our simulations.

### v2

@@ II Neutral atom error correction
Neutral atoms, such as 85Rb and 88Sr, often encode the qubit computational basis states in the electronic ground state manifold or clock states. These states have long coherence times on the order of seconds Jenkins et al. (2022); Ma et al. (2022) and the clock transition has a narrow linewidth Norcia et al. (2018). In the rest of this Article, we work in the $\{|0\rangle,|1\rangle,|r\rangle\}$ qutrit manifold of the 88Sr atom, which is endowed with the auxiliary Rydberg state $|r\rangle$ that facilitates long-range entanglement, which we leave unspecified. Fig. 1 shows the level scheme of 88Sr, in whose electronic states the $|0\rangle$- and $|1\rangle$-states have been embedded. The transition from $|1\rangle$ to $|r\rangle$ is driven by a single-photon process. Though we focus on ${}^{88}\text{Sr}$, our method applies to all neutral atoms with a similar level structure.
@@ II.1 Error modelling
Given is a noiseless unitary evolution $\mathcal{F}(\rho)=U\rho U^{\dagger}$. Introducing a noise channel with multi-qudit Kraus representatives $\{\mathfrak{D}_{\mu}\}$, we obtain a decoherent error channel
@@
relative to the perfect coherent channel $\mathcal{F}(\rho)$, for some coefficients $[\lambda_{\mu\nu}]$. These Kraus operators describe decoherent noise channels and satisfy the completeness relation
on the Hilbert space $\mathfrak{H}^{m}$ of $m$ qudits. The operators themselves are generated by single-qudit Kraus operators $E_{q_{i}}^{i}$ acting on qubit $i$, with index $q_{i}$. Through a change of basis, we can transform this channel $\mathcal{E}(\rho)$ to the Pauli basis as
@@
and use the Pauli twirling approximation Geller and Zhou (2013); Dankert et al. (2009); Jandura and Pupillo (2024) to obtain a channel which is diagonal in the $\{\mathcal{P}_{\mu}\}$-basis, by only keeping diagonal terms:
@@
The validity of twirling is discussed in Appendix A.
@@ II.2 Generating entanglement
Let $\mathbb{S}=\mathbb{S}_{X}\cup\mathbb{S}_{Z}$ be the group of stabilisers. Stabilisation between data qubits and stabiliser qubits $S\in\mathbb{S}$ is achieved through local 2-qubit entanglement operations. Neutral atoms provide a multitude of implementations, such as the simple constant pulses proposed by Jaksch et al. Jaksch et al. (2000) or a time-optimal pulse proposed by Jandura et al. Jandura and Pupillo (2022), which will be central to this Article. These are just a handful of examples of a very rich number of pulse protocols developed in the past few years, such as the pulse proposed by Pagano et al. Pagano et al. (2022) that achieves a pulse time very comparable to Ref. Jandura and Pupillo (2022) and has a lower integrated Rydberg time compared to the protocol introduced in Ref. Levine et al. (2019). Recently, also a pulse that is robust against time-dependent variations in the laser control parameters was proposed by Mohan et al. Mohan et al. (2023).
@@
Fig. 3(a) provides an implementation scheme of 5-qubit entanglement on the surface code. We implement subsequent 2-qubit gates in a clockwise fashion, and assume no interaction between Rydberg-excited data qubits. Movable tweezers, depicted in Fig. 3(b), can transport atoms within their stabilisers’ Rydberg blockade regime, where controlled logic can be performed, before moving them back to their initial positions in the lattice. Fig. 3(c) shows a complete schematic overview of how quantum error correction can be achieved on a neutral atom platform, showing an architecture with different dedicated zones for data storage, entanglement and readout.
@@
Figure 2: (a) The $[[9,1,3]]_{2}$ rotated surface code, where black circles denote data qubits that carry the 2 logical degrees of freedom, black faces denote $X$-stabilisers and golden faces denote $Z$-stabilisers. (b) For each stabiliser type, the relevant circuit is shown for 2-plaquettes. $H$ is the Hadamard gate and meas denotes a mid-circuit measurement. 4-plaquettes have analogous circuit designs.
@@
Figure 3: (a) Entanglement protocol for 4-plaquettes. The dashed circle indicates the Rydberg blockade radius, limited below the interatomic distance. Controlled logic between stabiliser and data qubits is performed by transporting the neighbouring qubits within the blockade radius, and applying an appropriate series of pulses before moving it back to its original position in the lattice. We subsequently entangle other data qubits in a clockwise fashion, as indicated by the grey arrows. (b) Visualisation of atom transport, with all relevant timescales indicated below Bluvstein et al. (2022); Tian et al. (2023); Covey et al. (2019). AOM (acousto-optical modulator) tweezers have the ability to pick up an atom from its stationary trap and move it to a different location. Such transport allows for flexibility during execution of an algorithm, though transpilation may pose a serious challenge Tan et al. (2024). (c) Schematic architecture for neutral atom quantum computing experiments, tailored here for quantum error correction. We embed the CSS code in a 2D plane of atoms (grey zone), here seen from the side. Single-qubit operations are performed in the grey zone, entanglement is mediated in the purple Rydberg zone, and measurements of the qubit states are performed in the green readout zone. AOMs transport atoms between layers. The Rydberg laser that mediates the transition $|1\rangle\leftrightarrow|r\rangle$ is global, as well as the laser that reads out qubit states through fluorescence. Hence, dedicated zones are required to avoid crosstalk. Single-qubit operations in the grey zone are local and can be tuned from lattice site to site, within realistic experimental bounds.
@@ III Statistical model of errors
Figure 4: (a) A CSS code is mapped to a spin gauge model with random bonds, with spins $s^{X}$ (black) and $s^{Z}$ (golden) representing stabilisers and bonds representing data qubits. Bonds between spins $s^{i}$ are ferromagnetic ($+J$) and favour long-range order, if the corresponding data qubit does not have an error that can be picked up by the stabiliser type $i$. Such bonds are depicted as blue dashed lines. (b) If a data qubit undergoes an error (red), triggering two flags (pink), its respective bond turns locally anti-ferromagnetic $(-J)$, indicated by a red dashed line.
@@ III.2 $\mathbb{Z}_{2}$ Gauge theory - Random plaquette model
Figure 5: (Top) Example of percolation on a 15$\times$15 spin system that is dual to the toric code, where spins are given by grey nodes. Bonds and vertices are randomly removed at a uniform rate $r=0.15$, leaving a subgraph on which top-bottom and left-right percolation can be found. Examples of logical support are indicated in green and black respectively. As periodic boundary conditions apply, percolation clusters must connect left-right and top-bottom on the same horizontal respectively vertical line. (Bottom) Percolation probability as a function of the fraction of unerased qubits, for system sizes $L=5,10,25$. Darker colours indicate larger lattices. The vertical black line indicates the threshold at around 0.75 ($r\approx 0.25$.
@@ IV Code performance analysis
The rest of this Section is dedicated to obtaining qubit fidelities for the toric code implemented on a neutral atom quantum device, assuming two major loss channels mediated by the fragile Rydberg state. We evaluate the local error channels for 4-plaquettes, extracting probabilities $\chi_{\mu\mu}$ for certain twirled Pauli error strings on the data qubits, which includes the probability $q$ of a false flag error on the stabilisers, and insert these into a 3D random plaquette gauge model. Additionally, erasure rates are incorporated, whose effects are a result of the integrated Rydberg lifetime of the qubits plus the finite lifetime of the tweezer traps. Because of the anisotropy of our system, spacelike and timelike Wilson loops have different sensitivities to variations in temperature. Our calculations have shown that disorder is more spacelike than timelike, so that we sample timelike Wilson loops only. The details of the Metropolis algorithm and handling phase transitions are presented in Appendix C Caracciolo et al. (1995); Palassini and Caracciolo (1999).
@@ IV.2 Handling erasure
Figure 9: (a) The phase diagrams for our quantum error correction implementation of the toric code, with stabiliser circuits being implemented by either the Jaksch or time-optimal protocol. The dark blue regions indicate the QEC regime where arbitrarily low logical qubit fidelities can be achieved, granted $d$ is sufficiently large. The light blue regions indicate regions where QEC is only viable if lost stabilisers are replenished from some fresh reservoir, alleviating some of the fidelity bounds. The blank regime entails all parameters $(\gamma,\bar{r})$ such that QEC is no longer possible. Error bars are a combined results of both standard Monte Carlo finite sampling errors for 2500 error distributions and 10.000 steps, and the uncertainty in $\gamma$ due to randomly distributing errors over a partially erased lattice yielding a different effective error rate standard deviation. Because Wilson loop calculations for $r\gtrapprox 0.35$ are very difficult to perform in a Monte Carlo setting, we use a fit to approximate the phase boundaries, using the knowledge that for the limiting case of erasure-only errors ($\gamma\to 0$), we must recover 2D percolation thresholds calculated in Sec. III.3. A set of realistic system parameters are highlighted in black dashed lines/grey regions as examples. (b) Typical graph of the behaviour of Wilson loop averages, normalised by perimeter/area. Each line is the average of 10 randomly chosen runs averaged over 10.000 steps each, at $r=0$ and $\gamma/\Omega=10^{-3}$, for both the ordered ($T<T_{c}$) and disordered ($T>T_{c}$) regime, displaying a clear transition from a perimeter to an area law. The left graph shows $-\ln{\mathcal{W}[\tau]/|\tau|}$ on the $y$-axis, while the right shows $-\ln{\mathcal{W}[\tau]/S_{\tau}}$. (c) Examples of when Wilson loops can be easily constructed. If the loop encloses no erased qubits, the enclosed area $S_{\tau}$ is equal to a Euclidean area. If some erasures are enclosed, then the area is effectively less. For larger $\bar{r}$, the calculation of the area/perimeter law for arbitrary loops becomes convoluted and imprecise.
@@ IV.3 Logical qubit fidelity and lifetime
$\gamma^{\text{Jaksch}}_{\text{th}}\approx(2.5\pm 0.2)\cdot 10^{-3}\Omega\quad\text{and}\quad\gamma^{\text{TO}}_{\text{th}}\approx(4.5\pm 0.2)\cdot 10^{-3}\Omega,$ (42)
which are consistent with literature on optimising neutral atom laser protocols for stabilisation for QEC Jandura and Pupillo (2024), as it provides an upper bound on error rate thresholds without requiring an (optimal) decoder. Note that the results agree very well because of the simplicity of the toric code, having local connectivity, and show compatibility with the range of state-of-the-art experimental parameters
@@
The errors in $\gamma$ from our simulations are of a higher order of magnitude than those produced by simulations such as provided by the stim Python package, though we stress that the rough order or magnitude of an error rate threshold is more important than a $\sim$5variance. For more complex topological codes with more intricate error models, however, these minor Monte Carlo errors may be acceptable trade-offs, as our model may achieve greater time efficiency compared to performing many rounds of decoding which can be slow and underestimate the maximum error rate threshold.
@@ Appendix A Twirling
For the Pauli twirling approximation, this corresponds to a physical insertion of Pauli operators, drawn at random on random locations within the circuit. It is not known exactly how accurate the twirling approximation is, but it has been employed often with excellent agreement with exact simulations of the full system. It is known that Pauli twirling reduces the density matrix to the space of 1-designs. Another option would be to invoke Clifford twirling, which forms a 3-design. However, the physical insertions of unitaries from the Clifford group $\mathcal{C}^{(2)}$ would cause our error propagation model to break down. These twirls are useful in efficient simulations nonetheless since it has been shown that random circuits with local connectivity are approximate $t$-designs and can be simulated classically using only polynomial time.
@@ Appendix B Calculating the entries of $[\chi_{\mu\nu}]$
On neutral atom quantum computers, measurement time is still the dominant timescale of a quantum error correction experiment, with $\tau_{\text{meas}}\approx 10$ ms being the order of magnitude for non-destructive fluorescence readout. Therefore, we can assume that during measurement, all remaining Rydberg states will have decayed back to the $|1\rangle$-states (or erased). Choosing $\bar{\delta t}\gg 1/\gamma$, and $\bar{\mathfrak{D}}_{\mu}=\mathfrak{D}_{\mu}(\bar{\delta t})$, the final $\Gamma$-matrix becomes
@@ Appendix C Monte Carlo details
where $\varphi(\cdot)$ is some universal dimensionless function, and $\nu$ is a critical exponent. Simulations were all performed on the GPU to boost efficiency, batching parallel runs by temperature points and error configurations. The plaquette gauge model has local gauge invariance with a local order parameter, so we can not employ FSS.
@@
We construct Wilson loops using square loops along the $xy$-planes of the model, since Pauli twirling shows there is more disorder along spacelike plaquettes than timelike plaquettes. We average over all Wilson loops that can be constructed, because of the isotropy of our model along spacelike plaquettes. In case of erasure, we only add loops to the thermodynamic averaging contribution that have no erased qubits on the loop, only inside. For $r\gtrapprox 0.15$, this signifies a major breakdown in our simulations. To ensure a good approximation to the partition function, all simulations are a result of 25000 equilibration steps and 25000 Monte Carlo samples.